### v2

@@ Appendix B Prompt Syntax
  • SubSequence : A sequence of amino acids or other chemical representations, such as SMILES, which may encompass a full sequence or a specific region. A SubSequence can begin with two special tokens: SUBMOLECULAR_ENTITY , followed by a token indicating the SubSequence type (e.g., CDR3_REGION ). These tokens are optional and may be omitted when only one SubSequence is present.
  • Molecule: A complete molecule, such as a protein chain or small molecule, which may contain multiple SubSequences corresponding to sub-regions within the molecule. Each Molecule is initiated with two special tokens - a general token indicating the entity’s hierarchical level, MOLECULAR_ENTITY_TOKEN, followed by a token specifying the molecule type (e.g., MOLECULAR_ENTITY_EPITOPE). Additionally, Molecules can be marked with natural start and end tokens to denote instances where truncation has occurred, either in the original database or due to sequence length constraints.
@@
In predictive tasks, relevant attribute values are masked, while in generative or masked language modeling (MLM) tasks, spans within SubSequences or entire SubSequences are masked. Tokens denoting entity types can also be masked for type prediction tasks. Additionally, each entity may possess alternative expressions or mutations, which can be employed for comparison tasks.
@@
Examples showing how a prompt is created for several downstream tasks can be found in Table LABEL:table_task_queries
@@ B.1 Modular Tokenizer and Meta Tokens
To support multiple modalities within a single prompt we have developed ”Modular Tokenizer” which allows to utilize different tokenizers within a single prompt by mapping tokens from different domains (like SMILES carbon ”C” and amino acid cysteine ”C”) to the same ID space.
We use ”Meta Tokens” of the format ⟨@TOKENIZER-TYPE=...⟩ to indicate that everything following this meta token, up to the next meta token (or the end of the prompt) should be tokenized with the defined tokenizer. For example, ⟨@TOKENIZER-TYPE=AA⟩ tokenizes amino-acids, while ⟨@TOKENIZER-TYPE=SMILES⟩ can tokenize SMILES. Since all of those ”sub tokenizers” must exist in a single vocabulary space, our modular tokenizer orchestrates that, and provides mechanism to avoid conflicts and for sub-tokenizers to co-exist. Additionally, we support additional instructions within a meta token beyond only expression which (sub) tokenizer should be used. For example, ⟨@TOKENIZER-TYPE=AA@MAX-LEN=1000⟩ allows to restrict the maximum length of the tokenized sequence, which provides more granular control compared to only controlling the overall total max sequence tokenized length. It is worth emphasizing - meta tokens, by themselves, do not get tokenized into any token. They serve as instructions for the modular tokenizer. Further details on the implementation can be found on https://github.com/BiomedSciAI/fuse-med-ml/tree/master/fuse/data/tokenizers/modular_tokenizer
@@ C.3 Datasets
CELLxGENE [26], a platform for single-cell transcriptomics data, was used to assemble a dataset of gene expression sequences from individual cells. After filtering for samples labeled as “cell” in the ’suspension_type’ field, we curated a dataset of 30 million samples. These samples were processed and converted into the Genformer format for use in model training.
@@ C.4 Multitask Pretraining Approach
ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained for seven tasks simultaneously using a multitask learning approach. Gradients from each task were aggregated before applying optimizer updates. This approach, combined with a custom query system, enables the model to learn from different tasks and domains during co-training.
@@ C.5 Pretraining Tasks
Language Model Tasks. Four language model tasks were defined: (1) amino-acid sequence representation of antibodies based on OAS database [23] (2) amino-acid sequence representation of general proteins based on Uniref90 [22, 71] (3) smiles representation of small molecules based on a mixture of PubChem [25] and Zinc databases [24] (4) Genformer format representations [41] of cell genes based on CELLxGENE [26]. In all language modeling tasks, we employed span-denoising (similar to T5 [20]) with a mean noise span length of 5 and a noise density of 0.15. Additionally, a special token was introduced per entity type to make the model aware of it (e.g., ⟨MOLECULAR_ENTITY_TYPE_ANTIBODY_HEAVY_CHAIN⟩). When sequences were available from different species, an additional special token was also introduced (e.g., ⟨ATTRIBUTE_ORGANISM_HUMAN⟩)
@@ Appendix D Additional Results
DTI ⟨@TOKENIZER-TYPE=AA⟩<MASK> ⟨@TOKENIZER-TYPE=AA⟩MOLECULAR_ENTITY⟩ ⟨MOLECULAR_ENTITY_GENERAL_PROTEIN CC=..⟨@TOKENIZER-TYPE=SMILES⟩MOLECULAR_ENTITY⟩ ⟨MOLECULAR_ENTITY_SMALL_MOLECULE AD..EOS ⟨@TOKENIZER-TYPE=SCALARS_LITERALS⟩ {standardized pKd}

### v3

@@ Appendix B Prompt Syntax
  • Sequence : A sequence of amino acids or other chemical representations, such as SMILES, which may encompass a full sequence or a specific region.
  • Molecule: A complete molecule, such as a protein chain or small molecule, which may contain multiple sequences corresponding to sub-regions within the molecule. Each molecule is initiated with two special tokens - a general token indicating the entity’s hierarchical level, MOLECULAR_ENTITY, followed by a token specifying the molecule type (e.g., MOLECULAR_ENTITY_EPITOPE). Additionally, molecules can be marked with natural start and end tokens to denote instances where truncation has occurred, either in the original database or due to sequence length constraints.
@@
To support a wide range of tasks, MAMMAL utilizes special placeholder tokens of the form ⟨SENTINEL_ID_?⟩. These tokens serve as positionally aware anchors in the decoder output and are designed to flexibly handle both structured and unstructured prediction targets. For complex generation tasks, such as sequence infilling, molecule editing, or antibody design, multiple sentinel tokens (e.g., ⟨SENTINEL_ID_0⟩, ⟨SENTINEL_ID_1⟩ allow the model to generate multiple, disjoint segments of output. This enables fine-grained control over where the model should insert or complete information.
In simpler settings such as classification or regression, a single sentinel token (e.g., ⟨SENTINEL_ID_0⟩ is used to indicate the position where the model should output the prediction. This unified approach allows all tasks, whether discrete or continuous, generative or discriminative, to be framed consistently as a form of sequence-to-sequence learning, while preserving flexibility across domains. The use of sentinel tokens thus enables seamless task specification and modular decoding, making the model architecture adaptable to a broad spectrum of biomedical problems.
@@
Examples showing how a prompt is created for several downstream tasks can be found in Table LABEL:table_task_queries.
@@ B.1 Modular Tokenizer and Meta Tokens
To support multiple modalities within a single prompt, we developed ”Modular Tokenizer” which allows to utilization of different tokenizers within a single prompt by mapping tokens from different domains (like SMILES carbon ”C” and amino acid cysteine ”C”) to the same ID space.
We use meta tokens of the format ⟨@TOKENIZER-TYPE=...⟩ to indicate that everything following this meta token, up to the next meta token (or the end of the prompt) should be tokenized with the defined tokenizer. For example, ⟨@TOKENIZER-TYPE=AA⟩ tokenizes amino-acids, while ⟨@TOKENIZER-TYPE=SMILES⟩ can tokenize SMILES. Since all of those ”sub tokenizers” must exist in a single vocabulary space, our modular tokenizer orchestrates that, and provides mechanism to avoid conflicts and allow for sub-tokenizers to co-exist. Additionally, we support additional instructions within a meta token, beyond just the expression of which (sub) tokenizer should be used. For example, ⟨@TOKENIZER-TYPE=AA@MAX-LEN=1000⟩ allows users to restrict the maximum length of the tokenized sequence, which provides more granular control compared to only controlling the overall total max sequence tokenized length. It is worth emphasizing that meta tokens, by themselves, do not get tokenized into any token. They serve as instructions for the modular tokenizer. Further details on the implementation can be found on https://github.com/BiomedSciAI/fuse-med-ml/tree/master/fuse/data/tokenizers/modular_tokenizer
@@ C.3 Datasets
CELLxGENE [89], a platform for single-cell transcriptomics data, was used to assemble a dataset of gene expression sequences from individual cells. After filtering for human samples labeled as “cell” in the ’suspension_type’ field, we curated a dataset of 30 million samples. The sequences for model training were created as ranked lists of gene names, similar to Geneformer [38], but with several differences. Instead of normalizing by median expression and ranking the now unique values of expression count, we log-normalized and binned the samples, followed by alphabetic sorting to make the task well-defined. The binning removes the ability of noisy reads to significantly affect the pretraining and alphabetic sorting ensures that for every masked token there is a unique correct answer.
@@ C.4 Multitask Pretraining Approach
ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained on seven tasks simultaneously using a multitask learning approach. Gradients from each task were aggregated before updating the optimizer. This approach, combined with a custom query system, enables the model to learn from different tasks and domains during co-training.
@@ C.5 Pretraining Tasks
Language Model Tasks. Four language model tasks were defined: (1) amino-acid sequence representation of antibodies based on OAS database [86]; (2) amino-acid sequence representation of general proteins based on Uniref90 [69, 93]; (3) SMILES representation of small molecules based on a mixture of PubChem [88] and Zinc databases [87]; and (4) Geneformer format representations [38] of cell genes based on CELLxGENE [89]. In all language modeling tasks, we employed span-denoising (similar to T5 [76]) with a mean noise span length of 5 and a noise density of 0.15. Additionally, a special token was introduced per entity type to make the model aware of it (e.g., ⟨MOLECULAR_ENTITY_TYPE_ANTIBODY_HEAVY_CHAIN⟩). When sequences were available from different species, an additional special token was also introduced (e.g., ⟨ATTRIBUTE_ORGANISM_HUMAN⟩)
@@ Appendix D Additional Results
DTI ⟨@TOKENIZER-TYPE=AA⟩<MASK> ⟨@TOKENIZER-TYPE=AA⟩MOLECULAR_ENTITY⟩ ⟨MOLECULAR_ENTITY_GENERAL_PROTEIN CC=…⟨@TOKENIZER-TYPE=SMILES⟩MOLECULAR_ENTITY⟩ ⟨MOLECULAR_ENTITY_SMALL_MOLECULE AD…EOS ⟨@TOKENIZER-TYPE=SCALARS_LITERALS⟩ {standardized pKd}
@@
Appendix E Additional results - AlphaFold3
Table S5: Performance comparison between MAMMAL (pre-trained and fine-tuned) and various AlphaFold3 (AF3) configurations for antibody-antigen binding prediction. The metrics are Area Under the Receiver Operating Characteristic curve (AUROC) and Area Under the Precision-Recall Curve (AUPRC)
Method AUROC AUPRC
MAMMAL Pre Trained 0.491 0.473
MAMMAL Fine Tuned 0.931 0.928
AF3 Heavy chain only ipTM 0.448 0.489
AF3 Heavy chain only pTM 0.496 0.514
AF3 Heavy+Light chains ipTM 0.504 0.519
AF3 Heavy+Light chains pTM 0.497 0.519
Control: AF3 H CD206 ipTM 0.455 0.5
Control: AF3 H CD206 pTM 0.66 0.583
Table S6: Performance comparison of nanobody-antigen binding prediction on three targets measured by AUPRC and AUROC, between MAMMAL (pre-trained and fine-tuned) and AF3.
Method AUPRC $\uparrow$ AUROC $\uparrow$
CD206 VWF TBG CD206 VWF TBG
MAMMAL Pre Trained 0.359 0.464 0.376 0.478 0.505 0.425
MAMMAL Fine Tuned 0.995 0.802 0.526 0.997 0.832 0.700
AF3 ipTM 0.434 0.358 0.953 0.565 0.402 0.976
AF3 pTM 0.526 0.318 0.942 0.651 0.552 0.964
Figure 6: (a) Evaluation of binding to an alternative HER2 representing ECD sequence and antibodies (heavy alone or with light chain). CD206 shown as control. pTM shown for completeness in addition to ipTM. (b) Testing with an alternative HER2 ECD (Uniprot P04626) (c) Testing with a pretrained version of MAMMAL which was not fine-tuned on an antigen-antibody binding task, on antigen-antibody binding prediction (d) Testing AF3 pTM scores on antigen-nanobody binding prediction (e) TBG and three binding and three non-binding VHHs (nanobodies) as predicted by AF3.